5x+71y=100------(1)

x+45y=63---------(2)

Answer by Alan3354(69443)

5x+71y=100------(1)
x+45y=63---------(2)
------------
5x+71y = 100 ------(1)
5x+225y= 315 ------(2) times 5
----------------------------------------- Subtract